# A Mapper algorithm with implicit intervals and its optimization

### v1

@@ 4.2 Noisy dataset
Two disjoint circles 6 DBSCAN(0.6,5) 0.005 200
Two unequal disjoint circles 6 DBSCAN(0.3,5) 0.001 300
Two intersecting circles 5 DBSCAN(0.6,5) 0.01 100
@@
Two intersecting circles with noise 6 DBSCAN(0.2,5) 0.002 300
3D human 8 DBSCAN(0.1,5) 0.0001 350
@@ 5 Application
We set the number of intervals to $n=15$, the learning rate to $\gamma=0.001$, and the number of epochs to $N=300$. We apply agglomerative clustering with a threshold of $3.13$ (Müllner, 2011). Additionally, we initialize the parameters by fitting a Gaussian mixture model to the projected data. The graph modes of the GMM soft Mapper algorihtm without and with optimation is given in Appendix C. Training loss of the GMM soft Mapper while optimizing prameter is given in the Appendix D. The resulting optimized Mapper graph mode offers valuable insights into Alzheimer’s disease progression. The $\chi^{2}$ test is used to determine if there is a statistically significant difference between the distribution of the identified subgroup and the rest of the population (Mannan and Meslow, 1984).
In brain area 36, we find a distinct branch which is associated with high Braak scores, as shown in the red dashed box in Figure 6 (a). The bar charts in Figure 6 (b) reveals distribution differences between this branch and the rest nodes. The $P-$value of the $\chi^{2}$ test is $0.0047$, less than $0.05$, indicating a significant difference in distribution. $73\%$ of patients in this unique group have severe Alzheimer’s disease, while the averaged ratio in the rest nodes is $37\%$. This indicates that the gene expression patterns in this cohort differ from the rest, and merit further investigation.
@@ 6 Discussion
In this work, we introduce a novel approach for flexibly constructing a Mapper graph based on a probability model. We develop implicit intervals for soft Mapper graphs using a Gaussian mixture model and a multinomial distribution. With a given number of intervals, our algorithm automatically assigns each data point to an interval based on the allocation probability. Then based on these implicit intervals, we derive the concept of the Mapper graph mode as a point estimation. Additionally, we design an optimization approach that enhances the topological structure of Mapper graphs by minimizing a specific loss function. This function considers both the likelihood of the projected data with respect to the GMM and the topological information of the Mapper graph mode. This optimization process is particularly suited for complex and noisy datasets, as the standard Mapper algorithm can be sensitive to noise and may fail in certain situations. Both simulation and application studies demonstrate its effectiveness in capturing the underlying topological structures. The application of our proposed algorithm to gene expression of brain area 36 from the MSBB successfully identifies a unique subgroup whose distribution of Braak scores significantly differs from the rest of the nodes.
It is worth noting that the multimodal nature of the objective function makes it challenging to find a globally optimal Mapper graph. Careful parameter tuning and domain expertise are essential for generating a high quality Mapper graph. In addtion, the mean persistence loss function used in this work is a simple representation of information on the extended persistence diagram. Numerous alternative approaches may be available, for instance, one could consider distinguishing signals from noise in a Mapper graph using confidence sets (Fasy et al., 2013). Finding a more robust persistence-specific loss function will be a one of the furtur work directions. Morevoer, as you may have noticed, the filter function is also important for constructing a meaningful Mapper graph, in this work, covers are directly constructed on filtered data first, as most literture does. Optimizing the filter function and intervals simultaneously would be a challenging but meanfully direction.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (12401383), the Shanghai Science and Technology Program (No. 21010502500), the startup fund of ShanghaiTech University, and the HPC Platform of ShanghaiTech University
Appendix A. Theoretical derivation of the mode of a mutilnomial distribution
@@
To compute the mode of a Mutilnomial distribution, i.e.,
@@ Appendix B. Samples from the optimized GMM soft Mapper of synthetic datasets
Figure 9: Eight samples from the optimized GMM soft Mapper of the two disjoint circles with different radius dataset.
Figure 10: Eight samples from the optimized GMM soft Mapper of the two intersecting circles with different radius dataset.
@@
Appendix C. The GMM soft Mapper modes for brain area 36
Figure 14: The graph mode of the GMM soft Mapper without optimization for brain area 36.
Figure 15: The graph mode of the GMM soft Mapper with optimization for brain area 36.
Appendix D. Loss figures
Figure 16: Training loss of the GMM soft Mapper while optimizing prameter for each dataset. (a). Dataset of the two disjoint circles. (b) Dataset of the two intersecting circles. (c) Dataset of the two circles with different radius. (d) Dateset of the two intersecting circles with different radius. (e) Dataset of the two intersecting circles with small noises. (f) Dataset of the two intersecting circles with large noises. (g) The 3D human dataset. (h) Brain area 36 dataset.

### v2

@@ 4.4 Quantitatively summary of synthetic datasets
Two disjoint circles 6 DBSCAN(0.3,5) 0.005 200
Two intersecting circles 5 DBSCAN(0.2,5) 0.01 300
Two unequal disjoint circles 5 DBSCAN(0.35,5) 0.001 300
@@
Two intersecting circles with small noise 6 DBSCAN(0.2,5) 0.002 250
Two intersecting circles with big noise 6 DBSCAN(0.2,5) 0.001 300
3D human 8 DBSCAN(0.1,5) 0.0001 200
Table 2: Quantitative comparison of the Mapper, D-Mapper, and Mapper graph mode across different datasets. Generally, the D-Mapper has a higher value of metric $SC_{adj}$ than the standard Mapper and the Mapper graph mode performs best concerning metric $SC_{adj}$. The D-Mapper* refer to the D-Mapper but with a larger number of intervals ($K=8$).
Datasets Methods SC_norm TSR SC_adj
Two disjoint circles Mapper 0.59 1 0.8
D-Mapper 0.68 1 0.84
Mapper mode 0.73 1 0.86
Two intersecting circles Mapper 0.58 1 0.79
D-Mapper 0.59 0.33 0.46
D-Mapper* 0.59 1 0.79
Mapper mode 0.65 1 0.83
3D human Mapper 0.51 1 0.76
D-Mapper 0.5 1 0.75
Mapper mode 0.44 1 0.72
Two unequal-sized disjoint circles Mapper 0.61 1 0.8
D-Mapper 0.63 1 0.82
Mapper mode 0.7 1 0.85
Two unequal-sized intersecting circles Mapper 0.64 1 0.82
D-Mapper 0.63 1 0.82
Mapper mode 0.68 1 0.84
Two intersecting circles with small noises Mapper 0.52 0.29 0.40
D-Mapper 0.55 0.33 0.44
Mapper mode 0.5 1 0.77
Two intersecting circles with big noises Mapper 0.43 0.25 0.34
D-Mapper 0.47 0.25 0.36
Mapper mode 0.6 0.48 0.54
@@ 5 Application
We set the number of (implicit) intervals to $K=15$, the learning rate to $\gamma=0.001$, and the number of epochs to $N=300$. We apply agglomerative clustering with a threshold of $3.13$ (Müllner, 2011). Additionally, we initialize the parameters by fitting a Gaussian mixture model to the projected data. The graph modes of the GMM soft Mapper algorihtm without and with optimation is given in Appendix D. Training loss of the GMM soft Mapper while optimizing prameter is given in the Appendix E. The resulting optimized Mapper graph mode offers valuable insights into Alzheimer’s disease progression. The $\chi^{2}$ test is used to determine if there is a statistically significant difference between the distribution of the identified subgroup and the rest of the population (Mannan and Meslow, 1984).
In brain area 36, we find a distinct branch which is associated with high Braak scores, as shown in the red dashed box in Figure 6 (a). The bar charts in Figure 6 (b) reveals distribution differences between this branch and the other nodes. The $P-$value of the $\chi^{2}$ test is $0.0047$, less than $0.05$, indicating a significant difference in distribution. $73\%$ of patients in this unique group have severe Alzheimer’s disease (with Braak scores $=5$ or $=6$), while the averaged ratio in the rest nodes is $36.8\%$. This indicates that the gene expression patterns in this cohort differ from the rest, and merit further investigation. We also run the standard Mapper and D-Mapper on this dataset and the results are shown in the Figure 17 in Appendix D. The standard Mapper exhibits a distinct branch (in the red dashed box) that demonstrates a significant distributional difference compared to other nodes ($81\%$ severe conditions versus $37.1\%$ severe conditions with the $P-$value of the $\chi^{2}$ test is $0.0024$), as indicated in Figure  17 (a). However, this branch has more nodes and sub-branches, which indicates a more separate structure, than that of the Mapper graph mode. No similar branch is observed in the D-Mapper graph and it contains many scattered nodes (or connected components). This also further supports that one of the advantages of our algorithm is its ability to produce a more concise topological structure of data.
@@ 6 Discussion
In this work, we introduce a novel approach for flexibly constructing a Mapper graph based on a probability model. We develop implicit intervals for soft Mapper graphs using a Gaussian mixture model and a multinomial distribution. With a given number of implicit intervals, our algorithm automatically assigns each data point to an implicit interval based on the allocation probability. Then based on these implicit intervals, we derive the concept of the Mapper graph mode as a point estimation. Additionally, we design an optimization approach that enhances the topological structure of Mapper graphs by minimizing a specific loss function. This function considers both the likelihood of the projected data with respect to the GMM and the topological information of the Mapper graph mode. This optimization process is particularly suited for complex and noisy datasets, as the standard Mapper algorithm can be sensitive to noise and may fail in certain situations. Both simulation and application studies demonstrate its effectiveness in capturing the underlying topological structures. The application of our proposed algorithm to gene expression of brain area 36 from the MSBB successfully identifies a unique subgroup whose distribution of Braak scores significantly differs from the rest of the nodes.
It is worth noting that the multimodal nature of the objective function makes it challenging to find a globally optimal Mapper graph. We suggest performing multiple optimization runs with careful parameter tuning and incorporating domain knowledge to generate high-quality Mapper graphs. In addtion, the mean persistence loss function used in this work is a simple representation of information on the extended persistence diagram. Numerous alternative approaches may be available, for instance, one could consider distinguishing signals from noise in a Mapper graph using confidence sets (Fasy et al., 2013). Finding a more robust persistence-specific loss function will be a one of the future work directions. Moreover, as you may have noticed, the filter function is also important for constructing a meaningful Mapper graph. In this work, covers are constructed based on the filtered data, as in most literature. Optimizing the filter function and (implicit) intervals simultaneously would be a challenging but meaningful direction. Moreover, a more appropriate quantitative assessment of the Mapper algorithms poses significant challenges. We adopt $SC_{adj}$ in this work, but this measure still has several limitations, for example, how to balance the trade-off between the clustering performance and the topological information. Besides, the favour of fine clustering of $SC$ may also limit the performance of the $SC_{adj}$. Developing a more robust and objective quantitative evaluation metric for Mapper algorithms constitutes a critical open problem in the field. Lastly, how to use a Mapper algorithm to achieve association studies between genetic sites and phenotypes simultaneously is an insightful direction, for example, which will further empower the applications of the Mapper algorithms in computational biology.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (12401383), the Shanghai Science and Technology Program (No. 21010502500), the startup fund of ShanghaiTech University, and the HPC Platform of ShanghaiTech University.
Appendix A. Theoretical derivation of the mode of a multinomial distribution
@@
To compute the mode of a multinomial distribution, i.e.,
@@ Appendix B. Samples from the optimized GMM soft Mapper of synthetic datasets
Figure 9: Eight samples from the optimized GMM soft Mapper of the two disjoint circles with different radii dataset.
Figure 10: Eight samples from the optimized GMM soft Mapper of the two intersecting circles with different radii dataset.
@@
Table 3: Statistical analysis of samples from GMM soft Mapper. TC is short for two circles. TIC is short for two intersecting circles.
Configuration Metric Mean 95% CI Median Mode
TC Avg Degrees 1.98 [1.97, 1.99] 2.00 2.00
Connected Components 1.98 [1.97, 1.99] 2.00 2.00
Loops 2.29 [2.22, 2.35] 2.00 2.00
TIC Avg Degrees 2.00 [1.99, 2.01] 2.00 2.00
Connected Components 1.01 [1.00, 1.01] 1.00 1.00
Loops 1.15 [1.11, 1.20] 1.0 1.00
3D Human Avg Degrees 3.27 [3.22, 3.23] 3.22 3.25
Connected Components 1.00 [1.00, 1.01] 1.00 1.00
Loops 94.22 [89.93, 98.50] 70.0 58.0
Unequal TC Avg Degrees 1.64 [1.63, 1.65] 1.71 1.75
Connected Components 1.91 [1.90, 1.93] 2.00 2.00
Loops 0.75 [0.71, 0.78] 1.00 1.00
Unequal TIC Avg Degrees 2.62 [2.61, 2.63] 2.60 2.67
Connected Components 1.01 [1.00, 1.01] 1.00 1.00
Loops 11.22 [10.80, 11.63] 10.00 10.0
Small Noise Avg Degrees 1.96 [1.95, 1.97] 2.00 2.00
Connected Components 1.18 [1.16, 1.20] 1.00 1.00
Loops 1.27 [1.20, 1.35] 1.00 1.00
Big Noise Avg Degrees 3.27 [3.25, 3.28] 3.25 3.33
Connected Components 2.37 [2.34, 2.40] 2.00 2.00
Loops 1296.32 [1216.40, 1376.24] 790.50 120
Appendix C. More explanations for the D-Mapper*
For the two intersecting circles dataset with the number of intervals $K=8$, the results are shown in Figure 14. The experiment shows that the D-Mapper achieves a higher $TSR$ with a larger $K$, it is likely because that the increased number of groups in a GMM leads to more fine components. However, we should be aware that a large number of intervals may also lead to spurious structures in the Mapper graph.
Figure 14: D-Mapper graph of two intersecting circles dataset with number of intervals $K=8$.
Appendix D. Supplementary figures for MSBB dataset of the brain area 36
Figure 15: The graph mode of the GMM soft Mapper without optimization for the MSBB dataset of the brain area 36.
Figure 16: The graph mode of the GMM soft Mapper with optimization for the MSBB dataset of the brain area 36.
Figure 17: (a) The standard Mapper graph for the MSBB dataset of the brain area 36. (b) The D-Mapper graph for the MSBB dataset of the brain area 36.
Appendix E. Loss figures
Figure 18: Training losses of the GMM soft Mapper while optimizing prameters across different datasets. (a) The dataset of the two disjoint circles. (b) The dataset of the two intersecting circles dataset. (c) The dataset of the circles with different radii dataset. (d) The dataset of the two intersecting circles with different radii dataset. (e) The dataset of the two intersecting circles with small noises. (f) The dataset of the two intersecting circles with large noises. (g) The 3D human dataset. (h) The MSBB dataset of the Brain area 36.